# New Multi-messenger Probe of Dark Matter-Nucleon Interactions from Ultra-high Energy Cosmic Ray Acceleration

### v1

@@ V Conclusion
Considering the recent spike in interest in spiky DM density profiles around supermassive black holes, it is very timely to find new observables to test the existence of such DM spikes. Conversely, it is also timely to ask what DM properties can be constrained in the presence of such spikes. In this work, we have introduced a novel method to constrain the DM-nucleon cross section within the spikes under various assumptions. We noted that relatively weak interactions between DM and SM baryons has the potential to disrupt the acceleration of ultra-high-energy CRs in dense environments near supermassive black holes. This is the case even after accounting for baryonic feedback that can suppress DM spikes.
@@
To set bounds on the values of these cross sections or make statements on the existence of DM spikes, we relied on a couple of assumptions. The first, and most crucial, is that CRs, particularly the heaviest elements like iron nuclei, are accelerated within DM spikes surrounding active galactic nuclei. If that is the case, they would fragment before reaching observed ultra-high energies, for sufficiently large $\sigma_{\chi p}$. By requiring that iron nuclei with energies $\sim 50$ EeV can be accelerated and escape without shattering, we derived conservative upper bounds on the DM–proton cross section. For a spiked DM profile in a Milky-Way-like galaxy, we obtained $\sigma_{\chi p}\lesssim 7\times 10^{-29}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$, improving to $\sigma_{\chi p}\lesssim 9\times 10^{-31}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ when accounting for acceleration region geometry. We also examined whether DM could induce a GZK-like cut-off for iron nuclei within the AGN itself. From this, we set a bound on the DM–proton scattering cross section of $\sigma_{\chi\gamma}\lesssim 9\times 10^{-31}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ assuming a spiked profile for MW-like galaxies and $\sigma_{\chi\gamma}\lesssim 1\times 10^{-33}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ for NGC 1068-like galaxies.
@@
These bounds probe DM interactions at momentum transfers $\gtrsim 10$ MeV, offering complementary and in some cases stronger constraints than those from direct detection or cosmological observations. In addition, given the large momentum transfers probed in this system, our bounds are also relevant to inelastic dark matter scenarios [Tucker-Smith:2001myb], where the interactions of DM particles with nucleons are necessarily inelastic, requiring a minimum amount of momentum exchange to produce the excited DM state, $\chi p\to\chi^{*}p$. These scenarios are much less constrained by direct detection experiments, though model-dependent limits from accelerators also apply [Izaguirre:2015zva, Berlin:2018jbm] (see also [Acevedo:2025rqu]). Our results demonstrate that ultra-high-energy CR survival and multi-messenger signatures from AGNs can provide powerful and model-independent constraints on DM. We emphasize that more robust and improved limits can be obtained by incorporating dark matter spikes in more detailed CR acceleration models and studying the energy dependence of the composition of CRs as observed at Earth.
@@ Acknowledgments
We are grateful for helpful discussions with Chris Cappiello, Dibya Chattopadhyay, Peter Cox, Raj Gandhi, Avirup Ghosh, Jayden Newstead and Yago Porto. This work was supported by the Australian Research Council through Discovery Project DP220101727 plus the University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. This work was partially supported by the University of Iowa’s Year 2 P3 Strategic Initiatives Program through funding received for the project entitled “High Impact Hiring Initiative (HIHI): A Program to Strategically Recruit and Retain Talented Faculty.” The work of B.D. was partly supported by the US Department of Energy under grant No. DE-SC0017987 and by a Humboldt Fellowship from the Alexander von Humboldt Foundation. S.A.M. and B.D. wish to acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP*) and the Institute for Underground Science at SURF for hospitality and for providing a stimulating environment, where a part of this work was done. B.D. also thanks the organizers of WHEPP 2025 at IIT, Hyderabad for local hospitality during the final stages of this work.
@@ Appendix A Is there a spike?
The existence of a DM spike is not well established in the literature, see for example [Ullio:2001fb, Merritt:2002vj, Merritt:2003eu]. Specifically, one of the mechanisms for the destruction of DM spikes are (galactic) mergers. During such mergers, the spike can be tidally disrupted reducing the density and consequently $\gamma_{\mathrm{sp}}$. On the other hand [Bertone:2002je] argue that specifically for the Milky Way, this destruction mechanism is unlikely, mainly due to the fact that the center of the Milky Way is believed to have formed 12 billion years ago in a single merger, without subsequent ones.
Another option for the suppression of the spike is stellar heating and dynamical friction [Merritt:2003qk]. While the DM spike is partially argued to exist due to the steeply rising stellar density near the GC, these stars could in turn lower the density via kinetic heating and capture. These processes would shift the expected $\gamma_{\mathrm{sp}}\sim 7/3$ to $3/2$. While this is above the value predicted by the NFW profile, it is below the adiabatic prediction. Of note, is that observationally, there are claims that $\gamma\sim 2.3$ [chan2024first], which would be closer to a purely adiabatic regime. There are also suggestions, that the growth could happen instantaneously, not adiabatically. In such a scenario, while there would be a spike, it would be ”blunted” compared to the adiabatic regime [Ullio:2001fb]. We note that even though stellar heating could suppress the spike [DelaTorreLuque:2024wfz], even in the maximal heating scenario, the DM density is still well above the predictions of a pure NFW profile.
The main problem with the spiked profile is the Core-Cusp problem. The prediction of the spike builds on a cuspy (NFW) profile. Some measurements seem to show that the DM density is more cored, which would significantly reduce the DM spike. However, there are no measurements below kpc-scale to distinguish a spike profile from a cored one. Additionally, the MW may be cored  [Ou:2023adg, Lim:2023lss, Baumgart:2025dov].

### v2

@@ V Conclusion
Considering the recent spike in interest in spiky dark matter density profiles around supermassive black holes, it is very timely to find new observables to test the existence of such DM spikes. Conversely, it is also timely to ask what DM properties can be constrained in the presence of such spikes. In this work, we have introduced a novel method to constrain the DM-nucleon cross section within the spikes under various assumptions. We noted that relatively weak interactions between DM and SM baryons has the potential to disrupt the acceleration of ultra-high-energy CRs in dense environments near SMBHs. This is the case even after accounting for baryonic feedback that can suppress DM spikes.
@@
To set bounds on the values of these cross sections or to make statements on the existence of DM spikes, we relied on a couple of reasonable assumptions. The first, and most crucial, is that CRs, particularly the heaviest elements like iron nuclei, are accelerated within DM spikes surrounding active galactic nuclei. If that is the case, they would fragment before reaching observed ultra-high energies, for sufficiently large $\sigma_{\chi p}$. By requiring that iron nuclei with energies $\sim 50$ EeV can be accelerated and escape without shattering, we derived conservative upper bounds on the DM–proton cross section. For a spiked DM profile in a MW-like galaxy, we obtained $\sigma_{\chi p}\lesssim 7\times 10^{-29}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$, improving to $\sigma_{\chi p}\lesssim 3\times 10^{-31}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ when using NGC 1068-like galaxies.
We also examined whether DM could induce a GZK-like cut-off for iron nuclei within the AGN itself. From this, we set a bound on the DM–proton scattering cross section of $\sigma_{\chi p}\lesssim 9\times 10^{-31}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ assuming a spiked profile for MW-like galaxies and $\sigma_{\chi p}\lesssim 1\times 10^{-33}(m_{\chi}/\mathrm{GeV})\mathrm{cm}^{2}$ for NGC 1068-like galaxies.
@@
These bounds probe DM interactions at momentum transfers $\gtrsim 10$ MeV, offering complementary and, in some cases, stronger constraints than those from direct detection or cosmological observations. In addition, given the large momentum transfers probed in this system, our bounds are also relevant to inelastic DM scenarios [163], where the interactions of DM particles with nucleons are necessarily inelastic, requiring a minimum amount of momentum exchange to produce the excited DM state, $\chi p\to\chi^{*}p$. These scenarios are much less constrained by direct detection experiments, though model-dependent limits from accelerators also apply [110, 46] (see also Ref. [17]). Our results demonstrate that ultra-high-energy CR survival and multi-messenger signatures from AGNs can provide powerful and model-independent constraints on DM. We emphasize that more robust and improved limits can be obtained by incorporating DM spikes in more detailed CR acceleration models and studying the energy dependence of the composition of CRs as observed at Earth.
@@ Acknowledgments
We are grateful for helpful discussions with Chris Cappiello, Dibya Chattopadhyay, Peter Cox, Damiano Fiorillo, Raj Gandhi, Avirup Ghosh, Jayden Newstead and Yago Porto. This work was supported by the Australian Research Council through Discovery Project DP220101727 plus the University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. This work was partially supported by the University of Iowa’s Year 2 P3 Strategic Initiatives Program through funding received for the project entitled “High Impact Hiring Initiative (HIHI): A Program to Strategically Recruit and Retain Talented Faculty.” The work of B.D. was partly supported by the US Department of Energy under grant No. DE-SC0017987 and by a Humboldt Fellowship from the Alexander von Humboldt Foundation. S.A.M. and B.D. wish to acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP*) and the Institute for Underground Science at SURF for hospitality and for providing a stimulating environment, where a part of this work was done. B.D. also thanks the organizers of WHEPP 2025 at IIT, Hyderabad for local hospitality (and for arranging hospital visits) during the final stages of this work.
@@ Appendix A Is there a spike?
The existence of a DM spike is not well-established in the literature; see for example, the arguments for and against it in Refs. [164, 127, 128]. Specifically, one of the mechanisms for the destruction of DM spikes are (galactic) mergers. During such mergers, the spike can be tidally disrupted reducing the density and consequently $\gamma_{\mathrm{sp}}$. On the other hand, Ref. [49] argues that specifically for the MW, this destruction mechanism is unlikely, mainly due to the fact that the center of the MW is believed to have formed 12 billion years ago in a single merger, without subsequent ones.
Another option for the suppression of the spike is stellar heating and dynamical friction [129]. While the DM spike is partially argued to exist due to the steeply rising stellar density near the galactic center, these stars could in turn lower the density via kinetic heating and capture. These processes would shift the expected $\gamma_{\mathrm{sp}}\sim 7/3$ to $3/2$. While this is above the value predicted by the NFW profile, it is below the adiabatic prediction. Of note, is that observationally, there are claims that $\gamma\sim 2.3$ [60], which would be closer to a purely adiabatic regime. There are also suggestions, that the growth could happen instantaneously, not adiabatically. In such a scenario, while there would be a spike, it would be “blunted” compared to the adiabatic regime [164]. We note that even though stellar heating could suppress the spike [67], even in the maximal heating scenario, the DM density is still well above the predictions of a pure NFW profile.
The main issue with the spiked profile is the Core-Cusp problem. The prediction of the spike builds on a cuspy (NFW) profile. Some measurements, most notably the recent Gaia data [135, 122, 42] seem to suggest that the DM density at the center of MW is more cored, which would significantly reduce the DM spike. However, there are no measurements nor simulations below kpc-scale to distinguish a spiked profile from a cored one. Moreover, even though the MW may not host a spiked DM profile, AGNs with ongoing accretion and adiabatic growth of the SMBH are more likely to do so.